f(x)=-2x+5
I need to find the x-intercept of the linear equation as a reduced fraction.
This is what I did:
0=-2x+5
0-5=-2x+5-5
-5=-2x
-5/-2=-2x/-2
-5/-2=x
This answer was wrong. What am I doing wrong?

Answer by Alan3354(69443)

f(x)=-2x+5
I need to find the x-intercept of the linear equation as a reduced fraction.
This is what I did:
0=-2x+5
0-5=-2x+5-5
-5=-2x
-5/-2=-2x/-2
-5/-2=x
This answer was wrong. What am I doing wrong?
==================
You didn't do anything wrong, but the answer is probably expected to be
x = 5/2 not -5/-2